# Dipole representation of composite fermions in graphene’s quantum Hall systems

### v3

@@ II.3 The effective Hamiltonian in dipole represenation in graphene
The effective Hamiltonian in this framework must reflect the dipole representation within $\mathcal{N}$-th LL. Additionally, it must preserve PH symmetry, meaning it remains invariant under the exchange of particles and holes. The Hamiltonian is carefully constructed to satisfy these conditions, and we impose a constraint that ensures the resulting Hamiltonian has a PH symmetric form:
@@
Here, $V^{(\mathcal{N})}(\vec{q})=\frac{2\pi e^{2}}{\epsilon|\vec{q}|}e^{-q^{2}/2}\left(F^{(\mathcal{N})}\right)^{2}$ represents the effective interaction, which takes into account states within the $\mathcal{N}$-th LL. The form factor mimics the LL characteristics. Moreover, it should be noted that $\upsilon~{}(\vec{q})=\frac{2\pi e^{2}}{\epsilon|\vec{q}|}$ defines the Fourier transform the Coulomb interaction potential.
@@
respectively. It can be concluded that the effective interaction can be represented as the Coulomb interaction $1/r$ combined with short-range functions to account for deviations at short distances. Thus, in addition to the Coulomb interaction, we will also utilize the Toke and Park interactions in the subsequent analysis. The coefficients $c_{i}$ ($i=0,1,2,3,4,5,6$) for $V_{\text{Toke}}$ and $a_{1},a_{2},\alpha_{1},\alpha_{2}$ for $V_{\text{Park}}$ are determined by matching the effective interactions $V_{\text{Toke}}(q)$ and $V_{\text{Park}}(q)$, operating in the LLL, to the first seven ($m=0,1,2,3,4,5,6$) and four ($m=0,1,2,3$) pseudopotentials of the Coulomb interaction $V_{2m-1}^{(n)}$, respectively, in the second LL ($n=1$) and the fourth LL ($n=3$). The Fourier transforms of the effective interactions are obtained as:
@@ III Dipole representation of composite fermions in graphene
Figure 2: Self-consistently calculated solutions for $\Delta(\vec{k})$ for different pairing channels $l$ in the second ($\mathcal{N}=1$) and fourth ($\mathcal{N}=3$) Landau levels (LLs). These results correspond to the Toke and Park interactions.
@@
In the context of the Coulomb interaction, it is energetically favorable for these correlation holes to move away from the electrons, allowing real holes to surround the electrons instead. We have previously shown that the density of these correlation holes is equal to the density of real holes, and that the size of the resulting dipoles is determined by a translation operator. As a result, at the Fermi level $k=k_{F}$, the electrons are positioned far from the holes, making this FLL state energetically unstable and prone to become a critical FLL state.
Breaking the symmetry between the left (L) and right (R) components of the system is equivalent to breaking the symmetry between particles and real holes. Therefore, states that break this symmetry are likely to be more energetically favorable. It is worth noting that at half- filling, most of the paired states that emerge (such as the Moore-Read Pfaffian, anti-Pfaffian, and f-wave states) spontaneously break this symmetry, with the PH-Pfaffian being an exception.
@@
The dipole representation of these Hamiltonians plays a crucial role in defining a single energy, which is essential for obtaining paired solutions. Additionally, they can also be interpreted as symmetry-breaking modifications of the Hamiltonian in Eq. (16), stabilizing the system into one of two paired states.
@@
Figure 3: Comparison of the total energy for various pairing solutions, expressed as $E_{\text{paired}}(l)-E_{\text{paired}}(l=0)$. The interaction strength is given by $V_{0}=2\pi e^{2}/l_{B}$.
@@
Figure 4: Comparing of total energies of the pairing solutions $E_{\text{paired}}$ and the regularized state $E_{\text{FLL}}$ in the four lowest LLs ($\mathcal{N}=0,1,2,3$) in the case of Coloumb, Toke, and Park interactions. Here, we calculated $E_{\text{paired}}$ for $l=-1$, which minimizes the energy value and is therefore the most energetically favorable. The interaction strength is defined by $V_{0}=4\pi e^{2}/l_{B}$.

### v4

# Dipole representation of composite fermions in graphene quantum Hall systems

@@ II.3 The effective Hamiltonian in dipole represenation in graphene
The effective Hamiltonian in this framework must reflect the dipole representation within $\mathcal{N}$-th LL. Additionally, it must preserve PH symmetry, meaning it remains invariant under the exchange of particles and holes. The Hamiltonian is carefully constructed to satisfy these conditions, and we impose a constraint that ensures the resulting Hamiltonian has a PH symmetric form [16]
@@
Here, $V^{(\mathcal{N})}(\vec{q})=\frac{2\pi e^{2}}{\epsilon|\vec{q}|}e^{-q^{2}/2}\left(F^{(\mathcal{N})}\right)^{2}$ represents the effective interaction, which takes into account states within the $\mathcal{N}$-th LL. The form factor mimics the LL characteristics. Moreover, it should be noted that $\upsilon~{}(\vec{q})=\frac{2\pi e^{2}}{\epsilon|\vec{q}|}$ defines the Fourier transform of the Coulomb interaction potential.
@@
respectively. It can be concluded that the effective interaction can be represented as the Coulomb interaction $1/r$ combined with short-range functions to account for deviations at short distances. Thus, in addition to the Coulomb interaction, we will also utilize the Toke and Park interactions in the subsequent analysis. The coefficients $c_{i}$ ($i=0,1,2,3,4,5,6$) for $V_{\text{Toke}}$ and $a_{1},a_{2},\alpha_{1},\alpha_{2}$ for $V_{\text{Park}}$ are determined by matching the effective interactions $V_{\text{Toke}}(q)$ and $V_{\text{Park}}(q)$, operating in the lowest LL (LLL), to the first seven ($m=0,1,2,3,4,5,6$) and four ($m=0,1,2,3$) pseudopotentials of the Coulomb interaction $V_{2m-1}^{(n)}$, respectively, in the second LL ($n=1$) and the fourth LL ($n=3$). The Fourier transforms of the effective interactions are obtained as:
@@ III Dipole representation of composite fermions in graphene
Figure 2: Self-consistently calculated solutions for $\Delta(\vec{k})$ for different pairing channels $l$ in the second ($\mathcal{N}=1$) (left) and fourth ($\mathcal{N}=3$) (right) Landau levels (LLs). These results correspond to the Toke and Park interactions.
@@
In the context of the Coulomb interaction, it is energetically favorable for these correlation holes to move away from the electrons, allowing real holes to surround the electrons instead. We have previously shown that the density of these correlation holes is equal to the density of real holes, and that the size of the resulting dipoles is determined by a translation operator. As a result, at the Fermi level $k=k_{F}$, the electrons are positioned far from the holes, making this FLL state energetically unstable and prone to become a critical FLL state [23].
Breaking the symmetry between the left (L) and right (R) components of the system is equivalent to breaking the symmetry between particles and real holes. Therefore, states that break this symmetry are likely to be more energetically favorable. It is worth noting that at half- filling, most of the paired states that emerge (such as the Moore-Read Pfaffian, anti-Pfaffian, and f-wave states) spontaneously break this symmetry.
@@
The dipole representation of these Hamiltonians plays a crucial role in defining a single energy, which is essential for obtaining paired solutions. Additionally, they can also be interpreted as symmetry-breaking modifications of the Hamiltonian in Eq. (16), stabilizing the system into one of two paired states [23].
@@
Figure 3: Comparison of the total energy for various pairing solutions, expressed as $E_{\text{paired}}(l)-E_{\text{paired}}(l=0)$ for Coulomb interactions (upper), and Toke and Park interactions (lower). The interaction strength is given by $V_{0}=2\pi e^{2}/l_{B}$.
@@
Figure 4: Comparing of total energies of the pairing solutions $E_{\text{paired}}$ and the regularized state $E_{\text{FLL}}$ in the four lowest LLs in the case of Coulomb interactions (upper), Toke, and Park interactions (lower). Here, we calculated $E_{\text{paired}}$ for $l=-1$, which minimizes the energy value and is therefore the most energetically favorable. The interaction strength is defined by $V_{0}=4\pi e^{2}/l_{B}$.